# Integrative Learning of Intensity Fluctuations of Quantum Dots under Excitation via a Tailored Mixture Hidden Markov Model

### v1

@@ 3.3 Likelihood Derivation and Estimation Criterion
Here we derive the likelihood function and discuss maximum likelihood estimation with the EM algorithm.
To proceed, we use $\pi_{k,\boldsymbol{s}_{i,1}}$ to denote the probability of the initial state when the $i$th QD is in the $k$th cluster, such that $\pi_{k,h}=Pr(s_{i,1,h}=1\mid z_{i,k}=1)$ and $\sum_{h=1}^{M}\pi_{k,h}=1$. Let $\boldsymbol{\pi}_{k}=(\pi_{k,1},\ldots,\pi_{k,M})^{\rm T}$, and $\boldsymbol{\pi}=(\boldsymbol{\pi}_{1},\ldots,\boldsymbol{\pi}_{K})^{\rm T}$. Let $\boldsymbol{\Theta}$ denote the set of all the unknown parameters, including the parameters for the state distributions $(\epsilon_{h}^{(0)},\epsilon_{h}^{(1)},a_{h},b_{h})_{h=1,\ldots,M}$, the parameters in the transition matrices $(\boldsymbol{\Pi}_{k})_{k=1,\ldots,K}$, and the parameters for the initial states $\boldsymbol{\pi}$.
Two key properties of the Markov chain are utilized. First, the sequence $(s_{i,1},\ldots,s_{i,T})$ forms a Markov chain if $s_{i,t+1}\perp(s_{i,1},\ldots,s_{i,t-1})\mid s_{i,t}$, meaning that the future is conditionally independent of the past given the present state. Second, the condition $x_{i,t}\perp(s_{i,1},\ldots,s_{i,t-1},x_{i,1},\ldots,x_{i,t-1})\mid s_{i,t}$ holds, meaning that the present observation depends only on the present state. With these properties and the inflated Beta distribution defined in (1), the complete-data likelihood of each dot $i$, given the intensity series $\boldsymbol{x_{i}}$, the state series $\boldsymbol{s_{i}}$, and the cluster information $\boldsymbol{z}_{i}$, is given by
$L(\boldsymbol{\Theta}\mid\boldsymbol{x}_{i},\boldsymbol{s}_{i},\boldsymbol{z}_{i})=\prod_{k=1}^{K}\left\{\delta_{k}\pi_{k,\boldsymbol{s}_{i,1}}f(x_{i,1}\mid\boldsymbol{s}_{i,1})\prod_{t=2}^{T}(\boldsymbol{\Pi}_{k})_{\boldsymbol{s}_{i,t-1},\boldsymbol{s}_{i,t}}f(x_{i,t}\mid\boldsymbol{s}_{i,t})\right\}^{I(z_{i,k}=1)},$ (2)
@@
3.4 Computational Algorithm & Model Selection
3.4.1 Computational Algorithms
We derive an EM algorithm for optimization, which executes an E-step and an M-step iteratively until convergence.
The procedures of the E-Step are shown as follows. First, we utilize the complete-data log-likelihood function in (2) and the multiple expectation structure to define the conditional expectation function of the complete-data log-likelihood (5),
@@ 3.4.1 Computational Algorithms
$\begin{split}\alpha_{i,t,k,h}&=Pr[(x_{i,1},\ldots,x_{i,t})^{T},s_{i,t,h}=1\mid z_{i,k}=1,\boldsymbol{\Theta}^{(j)}],\\ \beta_{i,t,k,h}&=Pr[(x_{i,t+1},\ldots,x_{i,T})^{T},s_{i,t,h}=1\mid z_{i,k}=1,\boldsymbol{\Theta}^{(j)}].\end{split}$
Borrowing the notations from Rabiner (1989), the quantities that need to be computed in the E-step is listed below:
@@
$\begin{split}\delta_{k}^{(j+1)}&=\frac{\sum_{i=1}^{N}\tau_{i,k}(\boldsymbol{\Theta}^{(j)})}{N},\\ (\boldsymbol{\Pi}_{k})_{p,q}^{(j+1)}&=\frac{\sum_{i=1}^{N}\tau_{i,k}(\boldsymbol{\Theta}^{(j)})\{\sum_{t=2}^{T}\xi_{i,t,k,p,q}(\boldsymbol{\Theta}^{(j)})}{\sum_{i=1}^{N}\tau_{i,k}(\boldsymbol{\Theta}^{(j)})\{\sum_{t=2}^{T}\sum_{q^{\prime}=1}^{M}\xi_{i,t,k,p,q^{\prime}}(\boldsymbol{\Theta}^{(j)})},\\ (\epsilon_{h}^{(r)})^{(j+1)}&=\frac{\sum_{i=1}^{N}\sum_{k=1}^{K}\tau_{i,k}(\boldsymbol{\Theta}^{(j)})\sum_{t=1}^{T}\gamma_{i,t,k,h}(\boldsymbol{\Theta}^{(j)})I(x_{i,t}=r)}{\sum_{i=1}^{N}\sum_{k=1}^{K}\tau_{i,k}(\boldsymbol{\Theta}^{(j)})\sum_{t=1}^{T}\gamma_{i,t,k,h}(\boldsymbol{\Theta}^{(j)})}\quad\text{for}r\in\{0,1\},\\ (a_{h}^{(j+1},b_{h}^{(j+1)})&=\operatorname*{arg\,max}_{(a_{h},b_{h})}\sum_{i=1}^{N}\sum_{k=1}^{K}\tau_{i,k}(\boldsymbol{\Theta}^{(j)})\{\sum_{t=1}^{T}\gamma_{i,t,k,h}(\boldsymbol{\Theta}^{(j)})\}.\end{split}$ (6)
@@
All the details of the proposed EM algorithm are given in Section A of the Supplementary Material.
3.5 Model Selection & Inference
For selecting the number of clusters and the number of states, several information criteria are available, including AIC (Akaike, 1974), BIC (Schwarz, 1978), and the Integrated Completed Likelihood (ICL) (Biernacki et al., 2000). However, our empirical study shows that no criterion consistently performs better than the others, though they all aim to balance model complexity and model goodness of fit. In practice, we advocate combining resulting from several information criteria, utilizing domain knowledge, and examining the clustering patterns with varying number of clusters to reach conclusions. To obtain the standard errors of the estimated parameters, we utilize the SEM algorithm in Meng and Rubin (1991). Details are provided in Section A.5 of the Supplementary Material.
@@ 4.1 Simulation Setup
We simulate $N$ samples from a MHMM-$\beta$ model with $M$ states and $K$ clusters. First, the QDs are randomly allocated to different clusters with assignment probabilities $\delta_{k},k\in\{1,\ldots,K\}$, where $\sum_{k=1}^{K}\delta_{k}=1$. For each dot in cluster $k$, we generate its hidden-state sequence as a Markov chain, with the transition matrix $\boldsymbol{\Pi}_{k}$, $k\in\{1,\ldots,K\}$. We then simulate the intensities given the hidden state sequence by sampling from the inflated-Beta distribution with parameters $(a_{h},b_{h},\epsilon_{h}^{(0)},\epsilon_{h}^{(1)})$, $h\in\{1,\ldots,M\}$.
Here we mainly present simulation settings that mimic the QD application. The total number of dots $N\in\{50,100\}$, and the number of states and clusters are fixed to be $3$, i.e., $M=K=3$. The cluster probabilities $(\delta_{1},\delta_{2},\delta_{3})$ have two scenarios: a relatively balanced partition $(0.3,0.3,0.4)$ and an unbalanced partition $(0.7,0.2,0.1)$. The transition matrices are given by
$\boldsymbol{\Pi}_{1}=\begin{bmatrix}0.50&0.25&0.25\\ 0.25&0.50&0.25\\ 0.25&0.25&0.50\end{bmatrix},\boldsymbol{\Pi}_{2}=\begin{bmatrix}0.94&0.05&0.01\\ 0.01&0.92&0.07\\ 0.01&0.05&0.94\end{bmatrix},\boldsymbol{\Pi}_{3}=\begin{bmatrix}0.84&0.12&0.04\\ 0.06&0.73&0.21\\ 0.02&0.17&0.81\end{bmatrix}$
The parameters for each state are also set as follows: $(a_{1},b_{1},\epsilon_{1}^{(0)},\epsilon_{1}^{(1)})=(2,4,0.1,0.1)$, $(a_{2},b_{2},\epsilon_{2}^{(0)},\epsilon_{2}^{(1)})=(8,4,0.05,0.05)$, and $(a_{3},b_{3},\epsilon_{3}^{(0)},\epsilon_{3}^{(1)})=(10,2,0.01,0.01)$.
@@ 4.2 Competing Methods
We compare our proposed mixture hidden Markov model with inflated Beta distribution, denoted by MHMM-$\beta$, to its closest competitor, the Gaussian mixture HMM model, denoted as MHMM-G. We also include an oracle procedure, denoted as MHMM-$\beta^{*}$, as a benchmark, where the clustering information is assumed known.
@@ 4.3 Evaluation Metrics
We report the estimation error of the inflated Beta distributions with $\mbox{Er}(\widehat{\boldsymbol{\theta}})=\sum_{h=1}^{M}\mbox{Er}(\widehat{\boldsymbol{\theta}}_{h})$, where $\mbox{Er}(\widehat{\boldsymbol{\theta}}_{h})=\|\boldsymbol{\widehat{\theta}}_{h}-\boldsymbol{\theta}_{h}\|_{2}$, with $\widehat{\boldsymbol{\theta}}_{h}=(\widehat{a}_{h},\widehat{b}_{h},\widehat{\epsilon}_{h}^{(0)},\widehat{\epsilon}_{h}^{(1)})^{\rm T}$ and $\boldsymbol{\theta}_{h}=(a_{h},b_{h},\epsilon_{h}^{(0)},\epsilon_{h}^{(1)})^{\rm T}$ representing the estimated and true parameters for state $h$. Similarly, the estimation error of the transition matrices is $\mbox{Er}(\widehat{\boldsymbol{\Pi}})=\sum_{k=1}^{K}\mbox{Er}(\widehat{\boldsymbol{\Pi}}_{k})$, where $\mbox{Er}(\widehat{\boldsymbol{\Pi}}_{k})=\|\boldsymbol{\widehat{\Pi}}_{k}-\boldsymbol{\Pi}_{k}\|_{F}$, with $\boldsymbol{\widehat{\Pi}}_{k}$ and $\boldsymbol{\Pi}_{k}$ representing the estimated and true transition matrix for cluster $k$, respectively. To measure the accuracy of retrieving the true cluster pattern of the dots, we report the percentage of correct clustering: $\mbox{CC}=\sum_{i=1}^{N}I(\widehat{\boldsymbol{z}}_{i}=\boldsymbol{z}_{i})/N$. The simulation under each setting is replicated 100 times and the results are averaged.

### v2

@@ 3.3 Likelihood Derivation and Estimation Criterion
Here we derive the data likelihood function and discuss the maximum likelihood estimation of the parameters. To proceed, we use $\pi_{k,\boldsymbol{s}_{i,1}}$ to denote the probability of the initial state when the $i$th QD is in the $k$th cluster, such that $\pi_{k,h}=Pr(s_{i,1,h}=1\mid z_{i,k}=1)$ and $\sum_{h=1}^{M}\pi_{k,h}=1$. Let $\boldsymbol{\pi}_{k}=(\pi_{k,1},\ldots,\pi_{k,M})^{\rm T}$, and $\boldsymbol{\pi}=(\boldsymbol{\pi}_{1},\ldots,\boldsymbol{\pi}_{K})^{\rm T}$. Let $\boldsymbol{\Theta}$ denote the set of all the unknown parameters, including the parameters for the state distributions $(\epsilon_{h}^{(0)},\epsilon_{h}^{(1)},a_{h},b_{h})_{h=1,\ldots,M}$, the parameters in the transition matrices $(\boldsymbol{\Pi}_{k})_{k=1,\ldots,K}$, and the parameters for the initial states $\boldsymbol{\pi}$.
Two key properties of the Markov chain are utilized. First, the sequence $(\boldsymbol{s}_{i,1},\ldots,\boldsymbol{s}_{i,T})$ forms a Markov chain if $\boldsymbol{s}_{i,t+1}\perp(\boldsymbol{s}_{i,1},\ldots,\boldsymbol{s}_{i,t-1})\mid\boldsymbol{s}_{i,t}$, meaning that the future is conditionally independent of the past given the present state. Second, the condition $x_{i,t}\perp(\boldsymbol{s}_{i,1},\ldots,\boldsymbol{s}_{i,t-1},x_{i,1},\ldots,x_{i,t-1})\mid\boldsymbol{s}_{i,t}$ holds, meaning that the present observation depends only on the present state. With these properties and the inflated Beta distribution defined in (1), the complete-data likelihood of each dot $i$, given the intensity series $\boldsymbol{x}_{i}$, the state series $\boldsymbol{s}_{i}$, and the cluster information $\boldsymbol{z}_{i}$, is given by
$L(\boldsymbol{\Theta}\mid\boldsymbol{x}_{i},\boldsymbol{s}_{i},\boldsymbol{z}_{i})\\ =\prod_{k=1}^{K}\left\{\delta_{k}\pi_{k,\boldsymbol{s}_{i,1}}f(x_{i,1}\mid\boldsymbol{s}_{i,1})\prod_{t=2}^{T}(\boldsymbol{\Pi}_{k})_{\boldsymbol{s}_{i,t-1},\boldsymbol{s}_{i,t}}f(x_{i,t}\mid\boldsymbol{s}_{i,t})\right\}^{I(z_{i,k}=1)},$ (2)
@@
3.4 Computational Algorithm, Model Selection, & Inference
3.4.1 Computational Algorithm
We derive a generalized EM algorithm for optimization, which executes an E-step and an M-step iteratively until convergence.
The procedures of the E-Step are shown as follows. First, we utilize the complete-data log-likelihood function in (2) and the multiple expectation structure to define the conditional expectation function of the complete-data log-likelihood as follows,
@@ 3.4.1 Computational Algorithm
$\begin{split}\alpha_{i,t,k,h}&=P[(x_{i,1},\ldots,x_{i,t})^{\rm T},s_{i,t,h}=1\mid z_{i,k}=1,\boldsymbol{\Theta}^{(j)}],\\ \beta_{i,t,k,h}&=P[(x_{i,t+1},\ldots,x_{i,T})^{\rm T},s_{i,t,h}=1\mid z_{i,k}=1,\boldsymbol{\Theta}^{(j)}].\end{split}$
The quantities that need to be computed in the E-step are listed below:
@@
$\begin{split}\delta_{k}^{(j+1)}&=\frac{\sum_{i=1}^{N}\tau_{i,k}(\boldsymbol{\Theta}^{(j)})}{N},\\ \boldsymbol{\pi}_{k,h}^{(j+1)}&=\frac{\sum_{i=1}^{N}\tau_{i,k}(\boldsymbol{\Theta}^{(j)})\gamma_{i,1,k,h}(\boldsymbol{\Theta}^{(j)})}{\sum_{i=1}^{N}\tau_{i,k}(\boldsymbol{\Theta}^{(j)})},\\ (\boldsymbol{\Pi}_{k})_{p,q}^{(j+1)}&=\frac{\sum_{i=1}^{N}\tau_{i,k}(\boldsymbol{\Theta}^{(j)})\{\sum_{t=2}^{T}\xi_{i,t,k,p,q}(\boldsymbol{\Theta}^{(j)})\}}{\sum_{i=1}^{N}\tau_{i,k}(\boldsymbol{\Theta}^{(j)})\{\sum_{t=2}^{T}\sum_{q^{\prime}=1}^{M}\xi_{i,t,k,p,q^{\prime}}(\boldsymbol{\Theta}^{(j)})\}},\\ (\epsilon_{h}^{(r)})^{(j+1)}&=\frac{\sum_{i=1}^{N}\sum_{k=1}^{K}\tau_{i,k}(\boldsymbol{\Theta}^{(j)})\sum_{t=1}^{T}\gamma_{i,t,k,h}(\boldsymbol{\Theta}^{(j)})I(x_{i,t}=r)}{\sum_{i=1}^{N}\sum_{k=1}^{K}\tau_{i,k}(\boldsymbol{\Theta}^{(j)})\sum_{t=1}^{T}\gamma_{i,t,k,h}(\boldsymbol{\Theta}^{(j)})}\quad\text{for}r\in\{0,1\},\\ (a_{h}^{(j+1)},b_{h}^{(j+1)})&=\operatorname*{arg\,max}_{(a_{h},b_{h})}\sum_{i=1}^{N}\sum_{k=1}^{K}\tau_{i,k}(\boldsymbol{\Theta}^{(j)})\{\sum_{t=1}^{T}\gamma_{i,t,k,h}(\boldsymbol{\Theta}^{(j)})\}.\end{split}$ (6)
@@
In the above we have mainly focused on the key structures of the algorithm; all the details are provided in Section A of the Supplementary Material.
3.4.2 Initial Values
Due to the non-convex nature of the problem, the EM algorithm may be sensitive to initial values. We consider two strategies, i.e., the initial values can either be randomly generated or constructed from single-dot analysis. For the latter, specifically, we fit a 0/1 inflated hidden Markov model for each QD, and then apply the K-means algorithm on their estimated transition matrices (with proper alignment) to find the initial clustering pattern of the samples. In practice, we find that combining the multiple random start strategy with the single-dot based initialization leads to the most stable results. In our numerical studies, unless otherwise noted, we fit the model with 10 sets of random initial values as well as the initial values from the single-dot analysis, and the final solution is the one that leads to the highest likelihood value.
3.4.3 Model Selection & Inference
To select the number of clusters and the number of states, several information criteria are available, including AIC (Akaike, 1974), BIC (Schwarz, 1978), and Integrated Completed Likelihood (ICL) (Biernacki et al., 2000). Our empirical study shows that these criteria perform well in general, and, as expected, AIC tends to favor a more complex model than BIC and ICL. In practice, we advocate combining the results from several information criteria, using domain knowledge, and examining the clustering patterns of varying numbers of clusters to reach a final conclusion. To obtain the standard errors of the estimated parameters, we utilize the SEM algorithm in Meng and Rubin (1991). Details are provided in Section B of the Supplementary Material.
@@ 4.1 Simulation Setup
We simulate data from the MHMM-$\beta$ model with $M$ states and $K$ clusters. First, the QDs are randomly allocated to different clusters with mixing probabilities $\delta_{k},k\in\{1,\ldots,K\}$, where $\sum_{k=1}^{K}\delta_{k}=1$. For each dot in cluster $k$, we generate its hidden-state sequence as a Markov chain, with the transition matrix $\boldsymbol{\Pi}_{k}$, $k\in\{1,\ldots,K\}$. We then simulate the intensities given the hidden state sequence by sampling from the 0/1 inflated-Beta distribution with parameters $(a_{h},b_{h},\epsilon_{h}^{(0)},\epsilon_{h}^{(1)})$, $h\in\{1,\ldots,M\}$.
Here, we mainly focus on simulation settings that mimic the QD application. The total number of dots is set to $N=100$, the number of states is set to $M=3$, and the number of clusters is set to $K=3$. We consider various lengths of the individual QD series, i.e., $T\in\{250,500,1000,2000\}$. The mixing probabilities have two scenarios: a relatively balanced partition with $(\delta_{1},\delta_{2},\delta_{3})=(0.3,0.3,0.4)$ and an unbalanced partition with $(\delta_{1},\delta_{2},\delta_{3})=(0.7,0.2,0.1)$.
We consider several scenarios for transition matrices and state distributions.
Scenario 1. In the first scenario, we closely mimic the estimated MHMM-$\beta$ model from the real data application. The true transition matrices are
$\boldsymbol{\Pi}_{1}=\begin{bmatrix}0.848&0.127&0.025\\ 0.060&0.794&0.146\\ 0.017&0.186&0.796\end{bmatrix},\boldsymbol{\Pi}_{2}=\begin{bmatrix}0.795&0.205&0.000\\ 0.000&0.994&0.006\\ 0.000&0.001&0.999\end{bmatrix},\boldsymbol{\Pi}_{3}=\begin{bmatrix}0.938&0.057&0.005\\ 0.013&0.924&0.063\\ 0.001&0.067&0.932\end{bmatrix},$
and the parameters for the three states are set as
$(a_{1},b_{1},\epsilon_{1}^{(0)},\epsilon_{1}^{(1)})$ $=(2.195,5.183,0.025,0.000),$
$(a_{2},b_{2},\epsilon_{2}^{(0)},\epsilon_{2}^{(1)})$ $=(10.077,6.805,0.000,0.001),$
$(a_{3},b_{3},\epsilon_{3}^{(0)},\epsilon_{3}^{(1)})$ $=(11.658,3.227,0.000,0.017).$
Scenario 2. In the second scenario, we modify the above settings to make the three clusters more distinct and the proportions of 0/1 inflation larger. The transition matrices are given by
$\boldsymbol{\Pi}_{1}=\begin{bmatrix}0.500&0.250&0.250\\ 0.250&0.500&0.250\\ 0.250&0.250&0.500\end{bmatrix},\boldsymbol{\Pi}_{2}=\begin{bmatrix}0.940&0.050&0.010\\ 0.010&0.920&0.070\\ 0.010&0.050&0.940\end{bmatrix},\boldsymbol{\Pi}_{3}=\begin{bmatrix}0.840&0.120&0.004\\ 0.060&0.730&0.210\\ 0.020&0.170&0.810\end{bmatrix},$
and the parameters for the three states are set as
$(a_{1},b_{1},\epsilon_{1}^{(0)},\epsilon_{1}^{(1)})$ $=(2.000,4.000,0.100,0.010),$
$(a_{2},b_{2},\epsilon_{2}^{(0)},\epsilon_{2}^{(1)})$ $=(8.000,4.000,0.050,0.050),$
$(a_{3},b_{3},\epsilon_{3}^{(0)},\epsilon_{3}^{(1)})$ $=(10.000,2.000,0.010,0.100).$
The two scenarios differ primarily in how distinct the clusters are. In Scenario 1, the long-run probabilities ($\boldsymbol{\pi}_{1}=(0.213,0.443,0.344)$, $\boldsymbol{\pi}_{2}=(0.000,0.143,0.857)$, $\boldsymbol{\pi}_{3}=(0.104,0.461,0.435)$) indicate that Clusters 1 and 3 transit among the three states with similar probabilities, making them more challenging to separate. In contrast, in Scenario 2, the long-run probabilities for all three clusters ($\boldsymbol{\pi}_{1}=(0.333,0.333,0.333)$, $\boldsymbol{\pi}_{2}=(0.143,0.385,0.473)$, $\boldsymbol{\pi}_{3}=(0.192,0.365,0.443)$) suggest comparatively sharper cluster distinctions.
@@ 4.2 Competing Methods
We compare our proposed MHMM-$\beta$ methods with several competing methods. Its closest competitor is the Gaussian mixture HMM model, denoted as MHMM-G, which jointly analyzes all the QDs but uses Gaussian state distributions.
Another set of competitors is based on single-dot analysis, including the Gaussian HMM model (HMM-G), and the 0/1 inflated-Beta HMM model (HMM-$\beta$). In order to perform clustering via single-dot analysis, the K-means algorithm is applied to the estimated individual transition matrices. The transition matrices within each cluster are estimated by exponentiating the average of the log-transformed individual transition matrices (McKinney et al., 2006).
Besides the above methods, we also include an oracle procedure, denoted as MHMM-$\beta^{*}$, as a benchmark, where the clustering information is assumed to be known and the true parameter values are used as initialization.
@@ 4.3 Evaluation Metrics
To compare model estimation, we report the error in estimating the means of the state distributions, i.e., $\mbox{Er}(\widehat{\mu})=\sqrt{\sum_{h=1}^{M}(\widehat{\mu}_{h}-\mu_{h})^{2}}$, the error in estimating the variances of the state distributions, i.e., $\mbox{Er}(\widehat{\sigma}^{2})=\sqrt{\sum_{h=1}^{M}(\widehat{\sigma}_{h}^{2}-\sigma_{h}^{2})^{2}}$, the error in estimating the mixing probabilities, i.e., $\mbox{Er}(\widehat{\delta})=\sqrt{\sum_{k=1}^{K}(\widehat{\delta}_{k}-\delta_{k})^{2}}$, and the error in estimating the transition matrices, i.e., $\mbox{Er}(\widehat{\boldsymbol{\Pi}})=\sum_{k=1}^{K}\|\boldsymbol{\widehat{\Pi}}_{k}-\boldsymbol{\Pi}_{k}\|_{F}$. Here $\widehat{\mu}_{h}$, $\widehat{\sigma}_{h}^{2}$, $\widehat{\delta}_{k}$, and $\boldsymbol{\widehat{\Pi}}_{k}$ are the estimates of their true counterparts, $\mu_{h}$, $\sigma_{h}^{2}$, $\delta_{k}$, and $\boldsymbol{\Pi}_{k}$, respectively. For MHMM-$\beta$ and MHMM-$\beta^{*}$, we also report the estimation error of the inflated-Beta distribution parameters, i.e., $\mbox{Er}(\widehat{\boldsymbol{\theta}})=\sqrt{\sum_{h=1}^{M}\|\boldsymbol{\widehat{\theta}}_{h}-\boldsymbol{\theta}_{h}\|_{2}^{2}}$, with $\widehat{\boldsymbol{\theta}}_{h}=(\widehat{a}_{h},\widehat{b}_{h},\widehat{\epsilon}_{h}^{(0)},\widehat{\epsilon}_{h}^{(1)})^{\rm T}$ and $\boldsymbol{\theta}_{h}=(a_{h},b_{h},\epsilon_{h}^{(0)},\epsilon_{h}^{(1)})^{\rm T}$ representing the estimated and true parameters, respectively.
To measure the accuracy of retrieving the true cluster pattern of the dots, we report the percentage of correct clustering: $\mbox{CC}=\sum_{i=1}^{N}I(\widehat{\boldsymbol{z}}_{i}=\boldsymbol{z}_{i})/N$.
The simulation under each setting is replicated 100 times and the results are averaged. The potential label switching problem is solved by the Hungarian Method in linear sum assignment problem (Papadimitriou and Steiglitz, 1982).